Jack Co. produces, and sells CPA review courses for individual use. The following information is provided:...

  1. Jack Co. produces, and sells CPA review courses for individual use. The following information is provided:
    1. Conversion cost was $360,000 and was four times the prime cost
    2. Direct materials used in production equaled $75,000
    3. Cost of goods manufactured was $415,000
    4. Beginning work in process is one-half the cost of ending work in process
    5. There are no beginning or ending inventories for direct materials
    6. Cost of goods sold was 90 percent of cost of goods manufactured
    7. Beginning finished goods inventory was $16,500

Requirements:

  1. Prepare the cost of goods manufactured for the current year
  2. Prepare the cost of goods sold for the current year
Answer #1
Jack Co.
Schedule of Cost of Goods Manufactured
Particulars
Direct materials used $        75,000
Direct Labor (Bal Fig) $        15,000
Prime Cost (360000/4) $        90,000
Add: Manufacturing Overheads (360000- direct labor cost) $      345,000
Total manufacuring costs $      435,000
Add: Beginning inventory of Work in Process $        20,000
Total Cost of Work in process $      455,000
Less: Ending inventory of Work in Process $        40,000
Cost of Goods manufactured (Given) $      415,000

Calculations:

Let cost of ending work in process = x
cost of Beginning work in process =1/2 of x    = 0.5x
Total manufacuring costs+ Beginning inventory of Work in Process - Ending inventory of Work in Process = Cost of Goods manufactured
435000    +    0.5x   - x = $      415,000
435000-415000 = -0.5x + x
20000 = 0.5x
x = 40000
cost of ending work in process (x) = 40000
cost of Beginning work in process (0.5x) = 20000
Cost of Goods sold :-
Beginning inventory of Finished Goods $     16,500
Add: Cost of Goods manufactured $   415,000
Cost of Goods available for sale $   431,500
Less: Ending inventory of Finished Goods (bal fig) $     58,000
Cost of Goods sold (90% of 415000) $ 373,500

Feel free to ask any clarification, if required. Kindly provide feedback by thumbs up, if satisfied. It will be highly appreciated.
Thank you.